OpenStudy (anonymous):

Simplify 7 square root of 28

OpenStudy (anonymous):

\[7\sqrt{28}\]

OpenStudy (anonymous):

Simplify 3 cubed root of 108 \[3\sqrt[3]{108}\]

OpenStudy (anonymous):

Thank you @ryan123345 can you help me with the other please <3

OpenStudy (ryan123345):

yup :p

OpenStudy (anonymous):

okay cause I didn't get that answer...

OpenStudy (anonymous):

\[6\sqrt[3]{4}\] thats what I got? am I right?

OpenStudy (ryan123345):

ohh i see how you want it done :p

OpenStudy (anonymous):

yeah.. I'm sorry. lol

OpenStudy (ryan123345):

youu end up getting 6 3sqrt 4 at the end but the you have to bring down the 3^3. you forgot about that

OpenStudy (anonymous):

so it would be 9\[\sqrt[3]{4}\]

OpenStudy (ryan123345):

3 x 3 x 3 is 27

OpenStudy (anonymous):

lol thats not one of my choices boo

OpenStudy (anonymous):

\[3\sqrt[3]{108}\]\[3\sqrt[3]{27}\sqrt[3]{4}\]\[3x3\sqrt[3]{4}\]\[9\sqrt[3]{4}\]

OpenStudy (anonymous):

Thank you @HawkCrimson you're so lovely :)

OpenStudy (anonymous):

\[\left( 5x-4 \right)\frac{ 1 }{ 2 }=6\] can you help me with this one? :)

OpenStudy (anonymous):

ur welcome :)

OpenStudy (anonymous):

multipy everything by 2 and make x the subject of the formulae, tell me know if u get stuck

OpenStudy (ryan123345):